OpenStudy (anonymous):

Please help! What is the equation of the line passing through (–3, –9) and parallel to the line y = x + 6 in standard form? x + y = –6 x + y = 6 x – y = –6 x – y = 6 <--- My guess

OpenStudy (anonymous):

ur guess is correct

OpenStudy (nurali):

y = x + 6.. has slope = 1 so the parallel line through (-3,-9) must also have slope = 1 y = x + b plug in the point and solve for b -9 = -3 + b b = -6 and the line is y = x - 6 and in standard form... x - y = 6

OpenStudy (nurali):

if correct please click best response.

OpenStudy (skullpatrol):